OpenStudy (anonymous):

Solve x^4 - 11x^2 + 18 = 0.

OpenStudy (mertsj):

Factor. Find two numbers whose product is + 18 and sum is -11

OpenStudy (asib1214):

follow my steps. x^4 - 11x^2 + 18 = 0 x^4 - 9x^2 - 2x^2 + 18 = 0 (x^4 - 9x^2) (-2x^2 + 18) = 0 x^2(x^2 - 9) -2(x^2 - 9) = 0 (x^2 - 9) (x^2 - 2) = 0 when you expand these brackets, you'll go back to your original equation.